# Insights from Large-Scale Communications: Initiation

### v1

@@ 3.1 SDSS Data
Mean number density ($\bar{n}_{g}$) ${6.87}\times 10^{-3}{\,h^{3}\,{\rm Mpc^{-3}}}$ ${5.70}\times 10^{-4}{\,h^{3}\,{\rm Mpc^{-3}}}$ ${4.13}\times 10^{-4}{\,h^{3}\,{\rm Mpc^{-3}}}$
Size of cubes ($L$) $150{\,h^{-1}\,{\rm Mpc}}$ $200{\,h^{-1}\,{\rm Mpc}}$ $250{\,h^{-1}\,{\rm Mpc}}$
@@
No. of available cubes $201$ $1081$ $922$
Weighted count per cube $23074\pm 2126$ $4643\pm 422$ $6584\pm 434$
Size of galaxy $13.19\pm 4.66{\,{\rm kpc}}$ $23.95\pm 7.42{\,{\rm kpc}}$ $24.05\pm 10.45{\,{\rm kpc}}$
$u-r$ colour $2.43\pm 0.72$ $3.78\pm 1.90$ $2.94\pm 2.03$
@@
We use data from the Sloan Digital Sky Survey (SDSS) to find the information content associated with galaxy clustering in the observable Universe. Galaxy samples from three 3 different redshift zones are prepared using data from the $18^{th}$ data release (almeida23) of SDSS. DR18 incorporates refined data from all the prior programs of SDSS. However, we use data accumulated through the BOSS and the Legacy programs in this work. A Structured Query is used to retrieve the required data from SDSS Casjobs  [^1] https://skyserver.sdss.org/CasJobs/. To start with, conditions on the equatorial coordinates of the galaxies to choose galaxies within the angular span $135^{\circ}\leq RA\leq 225^{\circ}$ and $-2^{\circ}\leq Dec\leq 63^{\circ}$, as it provides a uniform sky coverage with high completeness. The left panel of Figure 4 shows the sky coverage of both BOSS and Legacy; the rectangular region shows the chosen span. In the first step, we prepare 3 quasi-magnitude-limited samples by applying constraints over the redshift range and the (K-corrected + extinction-corrected) $r$-band absolute magnitude ( right panel of Figure 4 ). The three conditions applied for the respective samples are tabulated in Table 1. This minimizes the variation in number density in the samples, reducing the Malmquist bias, due to which fainter galaxies are difficult to identify at higher redshifts. Note that the BOSS data we use here includes targets from both LOWZ ($0.430.15<z<$ 0.43) and CMASS ($0.700.43<z<$ 0.70). On the other end, we have Legacy data with higher incompleteness on small redshifts ($z<0.02$). Also, fibre collision will affect the selection of galaxies in close vicinity. Hence, choosing an appropriate magnitude and redshift range alone does not provide a completely uniform sample selection. The radial number density of the combined (BOSS+Legacy) distribution still exhibits background fluctuations in addition to the intrinsic variation in number density ( resulting solely from the growth of inhomogeneities in the matter density field ). To decouple the two effects we identify the radial selection function $n(r)$ by applying a Gaussian kernel ( $5\sigma$). The smoothed radial number density field now considered the selection function, represents the residual Malmquist bias that has to be filtered out. Using this selection function we assign a weight $w(r)$ to each galaxy lying in a given radial bin $r$ to $r+\delta r$. If $\bar{n}_{g}$ is the maximum value of the $n(r)$ within the entire range of radial comoving distance then $w(r)=\frac{\bar{n}_{g}}{n(r)}$. One must take the sum over the weights instead of adding a unit count for each galaxy. Put differently, for each galaxy found at the comoving radius $r$, we are also losing $w(r)-1$ neighbouring galaxies due to the limitations of the survey. The top panels of Figure 5 show the radial number density profile of the three samples, before and after applying the selection function. The bottom three panels show galaxy distributions in slices, from each of the three quasi-magnitude-limited samples, projected on the $X$-$Y$ plane; each point represents a galaxy and the point sizes are scaled proportionally with the respective weights. Note that the three samples have different values of $\bar{n}_{g}$; hence, the values of the weights are specific to the samples. The comoving Cartesian coordinates of each galaxy in the sample are found using the spectroscopic redshifts and equatorial coordinates. After we have the three samples, we carve out cubic regions of $150{\,h^{-1}\,{\rm Mpc}}$, $200{\,h^{-1}\,{\rm Mpc}}$ and $250{\,h^{-1}\,{\rm Mpc}}$ from Sample 1, Sample 2 and Sample 3, respectively. we curve out as many cubes as possible by taking successive strides in all three directions. Any two adjacent cubes share $80\%$ the side length or almost $50\%$ of their volume. Relevant information about three samples is tabulated in Table 1. For each sample, we have shown the (average) apparent size and $u-r$ colour of the constituent galaxies. The size is determined from the product of the angular diameter distance and the Petrosian radius ( in radians ) of the circle containing $90\%$ of its total light.
@@ 3.2 Nbody Data
We generate a number of dark matter distributions by running a cosmological N-body simulation with GADGET4 springel21. The simulation is performed by populating $128^{3}$ dark-matter particles inside a $500{\,h^{-1}\,{\rm Mpc}}$ cube and tracking their evolution up to redshift zero. Ten realisations are generated and eight disjoint cubes are further extracted from each by dissecting them in each direction. We finally have $80$ cubes of size $250{\,h^{-1}\,{\rm Mpc}}$ for each of the redshifts that we have access to. To analyse the evolution of information content for $\Lambda CDM$ cosmology we analyse the distributions of dark-matter particles at redshifts 00, $0.5$, $1$, and $2$. For each of the $80$ cubes, we randomly select $100000$ dark-matter particles to retain them in the final samples.
@@ 3.3 Biased distributions in redshift-space
$\mathbf{s}=\mathbf{r}\,\left[1+\frac{\mathbf{v}\cdot\mathbf{r}}{a\,H(a)\,\,||\mathbf{r}||^{2}}\right],$ (35)
@@ 3.3.2 Biased distribution
Galaxies are formed after baryonic matter follows the pre-existing DM halos to select the density peaks for consolidation preferentially. So they are considered to be positively biased by and large. We use the N-body data to mimic a couple of biased galaxy distributions along with unbiased and negatively biased ones as well. To apply the bias on the redshift-space distribution of DM particles, we use the selection function (model 1) proposed by cole98. For a given cube of dark-matter particles mapped in redshift-space, We apply a Cloud-In-Cell (CIC) technique with a mesh of $N_{c}^{3}$ grids to measure the density contrast ($\delta$) on the grids. Thereafter, the z-score of density contrast, $\nu(x)=\frac{\delta(x)-\mu}{\sigma}$ is found for each of the particles, where $\sigma=\sqrt{<\delta^{2}>}$ and $\mu=\langle\delta\rangle=0$. We then apply trilinear interpolation to determine $\nu$ at the particle positions from the $8$ adjacent grids. Next, the selection probability
@@
is applied to have a density-wise biased selection of the particles from the entire cube. We also find the base probability of the particles $p_{0}(\nu)=\frac{(1+\nu\sigma)}{N_{c}^{3}}$; i.e. the probability of finding a randomly selected particle to have contrast $\nu$. After applying the selection function, the selection probability of a particle with scaled contrasts $\nu$ into the final set is $p_{f}(\nu)=p_{0}(\nu)\,p_{s}(\nu)$. Next, a random probability $p_{r}$ between 0 and $max\{p_{f}(\nu)\}$ is generated for each particle in the distribution. The particles which satisfy the criteria $p_{r}<p_{f}$ are selected to represent the galactic mass. We randomly select $100000$ particles, to get the final samples from each biased and unbiased distribution for further testing effects of linear RSD and biasing. The density contrast ($\delta$) is calculated ( by using CIC ), for each particle in the newly formed biased samples. Finally, linear bias is estimated by comparing their $\delta$ profile with the unbiased sample. We use different combinations of $A$ and $B$ and then check the linear bias for the distribution using the technique described in C. The parameters are tweaked to achieve the desired values of $b_{1}$ through trial and error.
Figure 6: The distribution of the different mock samples ($z=0.5$) described in subsection 3.3, before and after applying RSD and bias. Distribution for three different bias values is shown along with the original distribution and its counterpart in redshifted space. The variation in density contrast is represented using the variable colour palette.
The suitable values of the parameters $A$ and $B$ which provide the desired values of linear bias $b_{1}$ are enlisted in Table 2. Figure 6 helps compare the redshift-space biased and unbiased distributions with the original real-space N-body data. All distributions have the same number of particles.
Table 2: Specifications for the mock biased distributions from redshifted N-body cubes.
@@ 4.1 Information Spectrum and Information Sharing
In subsection 2.3, we discuss how to track the distribution of information gain in different frequencies through the information spectrum. On the other hand, the average information gain shared between two points separated by a certain distance in space is measured through the mutual gain, which is presented in subsection 2.4. At first, we take the 3 samples from SDSS and analyze the distribution of information in the observable Universe. We carry out the same analysis for simulated N-body data at different redshifts and further check the effect of linear RSD and bias by comparing SDSS results with redshift-space biased N-body mock distributions. For each case, we search for the scale of homogeneity where the mutual gain goes to zero.
4.1.1 Large-scale information sharing and scale of homogeneity
@@
In the top-left panel of Figure 7, the information spectrum is presented for three samples representative of three redshift zones. We notice a higher degree of information gain is associated with the larger wavelengths (small $k$) for all three samples. The amplitude gradually falls with increasing $k$ and finally reaches a plateau at the high-frequency end. We notice an abrupt change in the slope of the information spectrum near $0.06{\,h\,{\rm Mpc^{-1}}}$ for all three samples. This fall in spectral amplitude persists up to the nonlinear regime ($>0.78{\,h\,{\rm Mpc^{-1}}}$); after which, the consistent drop in information slows down and remains steady up to the smallest scales probed. This non-diminishing tail is the characteristic signature of information gain. Due to the presence of the exponential term, it contains all higher-order moments of density perturbations. Hence, it is capable of tracing the collective information offsets caused by all the multi-point correlations existing in that space of relevance. We find that Sample 2 has a nearly identical profile as Sample 3, but $\sim 7\%$ more information on higher frequencies. Sample 1 has a relatively lower information gain than Sample 2 and Sample 3 on all wavelengths. This is non-intuitive at first sight. On lower redshifts, the amplitude of information gain is expected to get higher with the enhancement in the definition of the structures. However, for Sample 1, we observe the information gain to be falling more rapidly in the large $k$ range. The presence of redshift-space distortions partly explains this fall. which is explained later in detail. However, the primary reason for this decline is that Sample 1 is a collection of a different class of galaxy than the other two samples. From Table 1, we find that galaxies in Sample 1 are much smaller in size compared to the galaxies in the other two samples. A relatively higher absolute magnitude and bluer $u-r$ color also indicate the presence of a younger population in the galaxy sample. This means that, unlike the other two samples, Sample 1 consists of a fair number of field and satellite galaxies that are not representative of a higher degree of clustering and hence cause less information gain.
@@
The bottom-left panel of Figure 7 shows the spectral index of information gain. If the information spectrum follows the relation $\bar{\cal S}\propto k^{n_{s}-1}$, then the spectral index of information gain is given as $n_{s}(k)=\frac{\Delta\ln{\bar{\cal S}}}{\Delta\ln k}+1$. For all the samples, we find $n_{s}$ lying between 00 and $1$, with a red-tilt that points toward an inside-out picture of information sharing. The spectral index starts with $n_{s}=1$ at the smallest $k$ we probe. This shows a scale-independent nature of the information spectrum on the largest scales or earliest epochs. It then gradually starts reducing, eventually reaching a turnaround point where $n_{s}$ finds its minimum value. This is where information gain has the maximum scale dependence. For Sample 2 and Sample 3, we can find this turnaround near $0.2{\,h\,{\rm Mpc^{-1}}}$. For Sample 1 however, we see this drop with a much lower value of the spectral index ($n_{s}\sim 0.1$) and in a relatively higher frequency range ($0.3-0.5{\,h\,{\rm Mpc^{-1}}}$) where $\bar{\cal S}$ becomes almost inversely proportional to $k$. In other words, bigger the structure, greater the information it contains.
@@
Figure 7: [Top-left]: Information spectrum as a function of wave numbers. [Bottom-left]: This Figure shows the spectral index $n_{s}$, at different values of $k$. [Top-right]: Information sharing as a function of length scales. The same thing is shown in the inset but in the log-scale. [Bottom-right]: Slope of information sharing. A zoomed-in view is presented in the inset.
@@
The top-right panel of Figure 7 shows the mutual gain as a function of length scale. The large information sharing found on small scales ($<10{\,h^{-1}\,{\rm Mpc}}$) advocates the coherence of the local structures. Two points in close vicinity share a similar fate regarding the evolution of their neighbourhood; hence, there is a maximum reduction in uncertainty on smaller scales, resulting in high information gain. As we go towards higher length scales information sharing falls $60-70\%$ up to $20{\,h^{-1}\,{\rm Mpc}}$. The sharp peaks that feature in the $\frac{\Delta^{2}{\cal G}^{M}}{\Delta r^{2}}$ vs $r$ plot (bottom right panel of Figure 7), correspond to the length-scale for the sharpest fall. As we can notice, this characteristic scale is slightly different for each sample and is connected to their mean intergalactic separation. Beyond this scale, the existing coherence gets rapidly weakened. In Sample 1, the average information sharing approaches zero at about $\sim 140{\,h^{-1}\,{\rm Mpc}}$. However, the mean ${\cal G}^{M}$ does not completely diminish for Samples 2 and 3 anywhere within the scope. Nevertheless, while examining the second derivatives, we observe that $\frac{\Delta^{2}{\cal G}^{M}}{\Delta r^{2}}$ for Sample 2 and 3 gradually approaches zero, exhibiting a steadily reducing amplitude. This implies that $\frac{\Delta{\cal G}^{M}}{\Delta r}$ is getting closer to a constant value. Now, from the ever-reducing trend for the ${\cal G}^{M}(r)$, at the largest available length scales, with a constant slope, will force the mutual gain eventually diminish at a certain length scale. This scale however is beyond the explored extent for these two samples. Note that galaxies separated by a distance beyond the scale of homogeneity cannot affect the reduction in uncertainty about their prospective evolutionary trajectories.

### v2

@@ 3.1 SDSS Data
We use data from the Sloan Digital Sky Survey (SDSS) to find the information content associated with galaxy clustering in the observable Universe. Galaxy samples from three 3 different redshift zones are prepared using data from the $18^{th}$ data release [79] of SDSS. DR18 incorporates refined data from all the prior programs of SDSS. However, we use data accumulated through the BOSS and the Legacy programs in this work. A Structured Query is used to retrieve the required data from SDSS Casjobs  [^1] https://skyserver.sdss.org/CasJobs/. To start with, conditions on the equatorial coordinates of the galaxies to choose galaxies within the angular span $135^{\circ}\leq RA\leq 225^{\circ}$ and $-2^{\circ}\leq Dec\leq 63^{\circ}$, as it provides a uniform sky coverage with high completeness. The left panel of Figure 4 shows the sky coverage of both BOSS and Legacy; the rectangular region shows the chosen span. In the first step, we prepare 3 quasi-magnitude-limited samples by applying constraints over the redshift range and the (K-corrected + extinction-corrected) $r$-band absolute magnitude ( right panel of Figure 4 ). The three conditions applied for the respective samples are tabulated in Table 1. This minimizes the variation in number density in the samples, reducing the Malmquist bias, due to which fainter galaxies are difficult to identify at higher redshifts.
@@
Number density ${6.87}\times 10^{-3}{\,h^{3}\,{\rm Mpc^{-3}}}$ ${5.70}\times 10^{-4}{\,h^{3}\,{\rm Mpc^{-3}}}$ ${4.13}\times 10^{-4}{\,h^{3}\,{\rm Mpc^{-3}}}$
Size of cubes $150{\,h^{-1}\,{\rm Mpc}}$ $200{\,h^{-1}\,{\rm Mpc}}$ $250{\,h^{-1}\,{\rm Mpc}}$
@@
Number of cubes $201$ $1081$ $922$
Galaxies per cube $23074\pm 2126$ $4643\pm 422$ $6584\pm 434$
Avgerage galaxy size $13.19\pm 4.66{\,{\rm kpc}}$ $23.95\pm 7.42{\,{\rm kpc}}$ $24.05\pm 10.45{\,{\rm kpc}}$
$u-r$ color $2.43\pm 0.72$ $3.78\pm 1.90$ $2.94\pm 2.03$
@@
Note that the BOSS data we use here includes targets from both LOWZ ($0.430.15<z<$ 0.43) and CMASS ($0.700.43<z<$ 0.70). On the other end, we have Legacy data with higher incompleteness on small redshifts ($z<0.02$). Also, fibre collision will affect the selection of galaxies in close vicinity. Hence, choosing an appropriate magnitude and redshift range alone does not provide a completely uniform sample selection. The radial number density of the combined (BOSS+Legacy) distribution still exhibits background fluctuations in addition to the intrinsic variation in number density ( resulting solely from the growth of inhomogeneities in the matter density field ). To decouple the two effects we identify the radial selection function $n(r)$ by applying a Gaussian kernel ( $5\sigma$). The smoothed radial number density field now considered the selection function, represents the residual Malmquist bias that has to be filtered out. Using this selection function we assign a weight $w(r)$ to each galaxy lying in a given radial bin $r$ to $r+\delta r$. If $\bar{n}_{g}$ is the maximum value of the $n(r)$ within the entire range of radial comoving distance then $w(r)=\frac{\bar{n}_{g}}{n(r)}$. One must take the sum over the weights instead of adding a unit count for each galaxy. Put differently, for each galaxy found at the comoving radius $r$, we are also losing $w(r)-1$ neighbouring galaxies due to the limitations of the survey. The top panels of Figure 5 show the radial number density profile of the three samples, before and after applying the selection function. The bottom three panels show galaxy distributions in slices, from each of the three quasi-magnitude-limited samples, projected on the $X$-$Y$ plane; each point represents a galaxy and the point sizes are scaled proportionally with the respective weights. Note that the three samples have different values of $\bar{n}_{g}$; hence, the values of the weights are specific to the samples. The comoving Cartesian coordinates of each galaxy in the sample are found using the spectroscopic redshifts and equatorial coordinates. After we have the three samples, we carve out cubic regions of $150{\,h^{-1}\,{\rm Mpc}}$, $200{\,h^{-1}\,{\rm Mpc}}$ and $250{\,h^{-1}\,{\rm Mpc}}$ from Sample 1, Sample 2 and Sample 3, respectively. we curve out as many cubes as possible by taking successive strides in all three directions. Any two adjacent cubes share $80\%$ the side length or almost $50\%$ of their volume. Relevant information about three samples is tabulated in Table 1. For each sample, we have shown the (average) apparent size and $u-r$ color of the constituent galaxies. The size is determined from the product of the angular diameter distance and the Petrosian radius ( in radians ) of the circle containing $90\%$ of its total light.
@@ 3.2 Nbody Data
We generate multiple realisations of dark matter distributions by running a cosmological N-body simulation with GADGET4 [80]. The simulation is performed by populating $128^{3}$ dark-matter particles inside a $500{\,h^{-1}\,{\rm Mpc}}$ cube and tracking their evolution up to redshift zero. Ten realisations are generated and eight disjoint cubes are further extracted from each by dissecting them in each direction. We finally have $80$ cubes of size $250{\,h^{-1}\,{\rm Mpc}}$ for each of the redshifts that we have access to. To analyse the evolution of information content for $\Lambda CDM$ cosmology we analyse the distributions of dark-matter particles at redshifts 00, $0.5$, $1$, and $2$. For each of the $80$ cubes, we randomly select $100000$ dark-matter particles to retain them in the final samples.
Figure 6: The distribution of the different mock samples ($z=0.5$) described in subsection 3.3, before and after applying RSD and bias. Distributions for three different bias values are shown along with the original distribution and its unbiased counterpart in redshifted space. The variation in density contrast is represented using the variable colour palette.
@@ 3.3 Biased distributions in redshift-space
$\mathbf{s}=\mathbf{r}\,\left[1+\frac{\mathbf{v}\cdot\mathbf{r}}{a\,H(a)\,\,||\mathbf{r}||^{2}}\right],$ (3.1)
@@ 3.3.2 Biased distribution
Galaxies are formed after baryonic matter follows the pre-existing DM halos to select the density peaks for consolidation preferentially. So they are considered to be positively biased by and large. We use the N-body data to mimic a couple of biased galaxy distributions along with unbiased and negatively biased ones as well. To apply the bias on the redshift-space distribution of DM particles, we use the selection function (model 1) proposed by [81]. For a given cube of dark-matter particles mapped in redshift-space, We apply a Cloud-In-Cell (CIC) technique with a mesh of $N_{c}^{3}$ grids to measure the density contrast ($\delta$) on the grids. Thereafter, the z-score of density contrast, $\nu(x)=\frac{\delta(x)-\mu}{\sigma}$ is found for each of the particles, where $\sigma=\sqrt{<\delta^{2}>}$ and $\mu=\langle\delta\rangle=0$. We then apply trilinear interpolation to determine $\nu$ at the particle positions from the $8$ adjacent grids. Next, the selection probability
@@
Table 2: Average bias for the $80$ mock biased distributions obtained from the redshifted N-body cubes.
@@
is applied to have a density-wise biased selection of the particles from the entire cube. We also find the base probability of the particles $p_{0}(\nu)=\frac{(1+\nu\sigma)}{N_{c}^{3}}$; i.e. the probability of finding a randomly selected particle to have contrast $\nu$. After applying the selection function, the selection probability of a particle with scaled contrasts $\nu$ into the final set is $p_{f}(\nu)=p_{0}(\nu)\,p_{s}(\nu)$. Next, a random probability $p_{r}$ between 0 and $max\{p_{f}(\nu)\}$ is generated for each particle in the distribution. The particles which satisfy the criteria $p_{r}<p_{f}$ are selected to represent the galactic mass. We randomly select $100000$ particles, to get the final samples from each biased and unbiased distribution for further testing effects of linear RSD and biasing. The density contrast ($\delta$) is calculated ( by using CIC ), for each particle in the newly formed biased samples. Finally, linear bias is estimated by comparing their $\delta$ profile with the unbiased sample. We use different combinations of $A$ and $B$ and then check the linear bias for the distribution using the technique described in A.3. The parameters are tweaked to achieve the desired values of $b_{1}$ through trial and error. The suitable values of the parameters $A$ and $B$ which provide the desired values of linear bias $b_{1}$ are enlisted in Table 2. Figure 6 helps compare the redshift-space biased and unbiased distributions with the original real-space N-body data. All distributions have the same number of particles.
@@ 4 Results
4.1 Evolution of Global Information entropy
In subsection 2.4 we discuss the method to measure the global information entropy at a given epoch for a log-normal density distribution and introduce three dimensionless quantities $\Gamma_{D_{+}}$, $\Gamma_{a}$ and $\Gamma_{H}$ to study the dynamics of the matter and space of the universe through the entropic gain. In the right-hand side of Figure 7 we show these three quantities as a function of redshift or scalefactor. In the top left panel of Figure 7 we show the evolution of the density distribution responsible for causing the increment in the global entropic gain. The density distribution estimated for the $250{\,h^{-1}\,{\rm Mpc}}$ ( with voxel size $\sim 8{\,h^{-1}\,{\rm Mpc}}$ ) cube for the Nbody dark matter distribution is shown alongside the log-normal distribution ($z=0$) for comparison. In the bottom left panel of Figure 7 we show the global entropic gain as a function of the scale factor. A negative entropic gain or entropic drop indicates the dominance of underdense regions in the governance of the large-scale configuration of the Universe. The nonlinear growth in the entropic drop illustrates the increasing growth of inhomogeneities in the density field. However, the net effect of entropic gain up to $z=0$ is minuscule, i.e. $<1\%$ of the total information entropy of the Universe. On the top-right panel, we show $\Gamma_{D_{+}}$ as a function of scalefactor. We find that throughout all redshifts the growth rate of the Universe scales as $\sim\sqrt{{\cal G}^{U}}$. On the right-middle panel of Figure 7 we show the evolution of $\Gamma_{a}$. We find that the size of the Universe scales as $\sim\sqrt{{\cal G}^{U}}$ on large redshifts. The dependence of the expansion of space on the entropic drop gradually increases with time and on smaller redshifts the size of the Universe approaches a scaling of $\sim{\cal G}^{U}$. In the bottom-right panel, we show the variation of $\Gamma_{H}$ with scale factor. We note that the comoving expansion rate reduces with the increment in the entropic drop. An implication of this can be the increasing expansion rate which may result from the natural tendency of the cosmos to try and override the entropic drop. On larger redshifts, we get $\Gamma_{H}\sim{-0.75}$. As the Universe transitions from decelerated to accelerated expansion, $\Gamma_{H}$ has the maximum rate of change; it reduces again on smaller redshifts. The Hubble flow is comparatively less sensitive to the entropic drop at present with $\Gamma_{H}\sim{-0.45}$. Overall, the evolution of $\Gamma_{D_{+}}$, $\Gamma_{a}$ and $\Gamma_{H}$ for the N-body distribution is found to be almost identical to the log-normal distribution, which suggests that the change in these quantities are solely dependent on the choice of the cosmological model.
Figure 7: [Top-left]: Evolution of the density field. [Bottom-left]: Entropic gain as a function of scalefactor. [Top-right]: $\Gamma_{D_{+}}$ as a function of scalefactor. [Right-middle]: $\Gamma_{a}$ as a function of scalefactor. [Bottom-right]: $\Gamma_{H}$ as a function of scalefactor.
@@ 4.2 Information spectrum and Information Sharing
In subsection 2.5, we discuss how to track the distribution of specific information gain in different frequencies through the information spectrum. On the other hand, we present the mutual gain to measure the average information gain shared between two points separated by a certain distance in space. At first, we take the 3 samples from SDSS and analyze the distribution of information in the observable Universe. We carry out the same analysis for simulated N-body data at different redshifts and further check the effect of linear RSD and bias by comparing SDSS results with redshift-space biased N-body mock distributions. For each case, we search for the scale of homogeneity where the mutual gain goes to zero.
4.2.1 Information spectrum in SDSS
@@
In the top-left panel of Figure 8, the spherically averaged information spectrum is presented for three samples representative of three redshift zones. We notice a higher degree of information gain is associated with the larger wavelengths (small $k$) for all three samples. The amplitude gradually falls with increasing $k$ and finally reaches a plateau at the high-frequency end. We notice an abrupt change in the slope of the information spectrum near $0.05{\,h\,{\rm Mpc^{-1}}}$ for all three samples. This fall in spectral amplitude persists up to the nonlinear regime ($>0.8{\,h\,{\rm Mpc^{-1}}}$); after which, the consistent drop in information slows down and remains steady up to the smallest scales probed. This non-diminishing tail is the characteristic signature of the information spectrum. Due to the presence of the exponential term, it contains all higher-order moments of density perturbations. Hence, it is capable of tracing the collective information offsets caused by all the multi-point correlations existing in that space of relevance. However, it is not capable of extracting the information on phase differences between the different modes. In Figure 8 (top-left panel) we find an overall trend of the information spectrum decreasing with increasing redshift. Sample 2 has a nearly identical profile as Sample 3 but a higher amplitude throughout. Similarly, Sample 1 has a higher amplitude than both Sample 2 and 3 up to $k=0.2{\,h\,{\rm Mpc^{-1}}}$. Beyond $k=0.2{\,h\,{\rm Mpc^{-1}}}$ we notice a rapid fall of the spectral amplitude for Sample 1 which is non-intuitive at first sight. On lower redshifts, the amplitude of information gain is expected to get higher with the enhancement in the definition of the structures. However, for Sample 1, we observe the information gain to be falling more rapidly in the large $k$ range. The presence of redshift-space distortions partly explains this fall, which is discussed later on in detail. That said, Sample 1 is a collection of galaxies from a different class than the other two samples, which could be a significant contributing factor to this drop in information gain on higher frequencies. From Table 1, we find that galaxies in Sample 1 are much smaller in size compared to the galaxies in the other two samples. A relatively higher absolute magnitude and bluer $u-r$ color also indicate the presence of a younger population in the galaxy sample. This means that, unlike the other two samples, Sample 1 consists of a fair number of field and satellite galaxies that are not representative of a higher degree of clustering and are more uniformly distributed on small scales, compared to galaxies in Sample 2 and 3.
@@
The bottom-left panel of Figure 8 shows the spectral index of information gain. If the information spectrum follows the relation $\bar{\cal S}\propto k^{n_{s}-1}$, then the spectral index of information gain is given as $n_{s}(k)=\frac{\Delta\ln{\bar{\cal S}}}{\Delta\ln k}+1$. For all the samples, we find $n_{s}$ lying between 00 and $1$, with a red-tilt that points toward an inside-out picture of information sharing. The spectral index starts with $n_{s}=1$ at the smallest $k$ we probe. This shows a scale-independent nature of the information spectrum on the largest scales or earliest epochs. It then gradually starts reducing, eventually reaching a turnaround point where $n_{s}$ finds its minimum value. This is where the information gain has the maximum scale dependence. For Sample 2 and Sample 3, we can find this turnaround near $0.2{\,h\,{\rm Mpc^{-1}}}$. For Sample 1 however, we see this drop with a much lower value of the spectral index ($n_{s}\sim 0.1$) and in a relatively higher frequency range ($0.3-0.5{\,h\,{\rm Mpc^{-1}}}$) where $\bar{\cal S}$ becomes almost inversely proportional to $k$. In other words, bigger the structure, greater the information it contains.
@@
Figure 8: [Top-left]: Information spectrum as a function of wave numbers. [Bottom-left]: This figure shows the spectral index $n_{s}$, at different values of $k$. [Top-right]: Mutual gain as a function of length scales. The same thing is shown in the inset but in the log scale. [Bottom-right]: $2^{nd}$ derivative of mutual gain with length scale; a zoomed-in view is presented in the inset.
@@
The top-right panel of Figure 8 shows the mutual gain as a function of length scale. The large information sharing found on small scales ($<10{\,h^{-1}\,{\rm Mpc}}$) advocates the coherence of the local structures. Two points in close vicinity share a similar fate regarding the evolution of their neighbourhood; hence, there is a maximum reduction in uncertainty on smaller scales, resulting in high entropic gain. As we go towards higher length scales information sharing falls up to $60\%$ between 10 $10\,-\,20{\,h^{-1}\,{\rm Mpc}}$. The sharp peaks that feature in the $\frac{\Delta^{2}{\cal G}^{M}}{\Delta r^{2}}$ vs $r$ plot (bottom right panel of Figure 8), correspond to the length-scale for the sharpest fall. As we can notice, this characteristic scale is slightly different for each sample and is connected to their mean intergalactic separation. Beyond this scale, the existing coherence gets rapidly weakened. In Sample 1, the average information sharing goes to zero at $\sim 135{\,h^{-1}\,{\rm Mpc}}$. However, the mean ${\cal G}^{M}$ does not completely diminish for Samples 2 and 3 anywhere within the scope. Nevertheless, while examining the second derivatives, we observe that $\frac{\Delta^{2}{\cal G}^{M}}{\Delta r^{2}}$ for Sample 2 and 3 gradually approaches zero, exhibiting a steadily reducing amplitude. This implies that $\frac{\Delta{\cal G}^{M}}{\Delta r}$ is getting closer to a constant value. Now, from the ever-reducing trend for the ${\cal G}^{M}(r)$, at the largest available length scales, with a constant slope, will force the mutual gain eventually diminish at a certain length scale. This scale however is beyond the explored extent for these two samples. Note that galaxies separated by a distance beyond the scale of homogeneity cannot affect the reduction in uncertainty about their prospective evolutionary trajectories through gravitational interactions, but that does not imply causal disassociation